Dublin University Rifle Club

Version Date : April 23, 2002

DURC Competition Attendance Report

UCD Air Rifle Open 2

Distribution

Dr. Trevor West DUCAC Chairman
Aubrey Storey DUCAC Secretary
Terry McAuley DURC President and Director of Sport
Michelle Bennett Physical Recreation Officer
Tony O’Donnell Student Record
Susan Thompson Trinity News

Summary

Venue UCD Rifle Club, Belfield
Date 23-24 February 2002
Discipline 10m Air Rifle

DURC Attendees

Steven Watterson  Mark Dennehy
Ben Jones  Nithin Santhosh
Rob McAdoo  Ljiljana Skrba
Kate Fanning  Victoria McDowell

Scores

Mark Dennehy 541
Ben Jones 504
Nithin Santhosh 504
Kate Fanning 442
Victoria McDowell 413
Ljiljana Skrba 382

Prizes Won

Ben Jones Second Place, class C

Report

DURC took part in UCD’s second air rifle open of the year this weekend. Our 40th anniversary celebration took place the night before, and unfortunately all this celebrating meant that our numbers were depleted, as many people were still asleep when thy should be shooting!

Many congratulations to Nithin Santhosh, who shot a personal best of 504 and earns promotion to class C, breaking the 500 barrier for the first time. Ben Jones equalled Nithin’s score (after two successful challanges when shots were mis-scored). Mark Dennehy, though disapointed with his score, reached the final and moved himself up to sixth place when John Woodward of WTSC had a disasterous first shot. Steven Watterson was unable to shoot as supervisory duty claimed the first day of competition and the DURC anniversary claimed the second - a shame, as given Steven’s recent scores and those put in on the weekend, we probably missed out on second place in the overall competition.

Kate Fanning, Victoria McDowell and Ljiljana Skrba all shot well and improved on their personal best scores. All this means we’re going to have an excellent team for the colours match against UCD, which is taking place next term on April 27th.

Misé le meas,

Mark Dennehy, DURC Secretary